OpenStudy (ari67):

Eight less than 1/3 a number n is -13 as an equation

jhonyy9 (jhonyy9):

8 < (1/3)n -13 hence is right ?

jhonyy9 (jhonyy9):

multiplie both sides by 3 24 < n-39 add to both sides 39 24+39 < n n > 63 hope helped

OpenStudy (sshayer):

\[\frac{ 1 }{ 3 }n-8=-13\] multiply by 3 n-24=-39 n=-39+24=-15 n=-15

OpenStudy (sshayer):

no need to solve only first line is required in your question.